OpenStudy (anonymous):

if you are given a triangle with lengths of 3 inches, 6 inches and 7 inches, is it a right triangle?

OpenStudy (anonymous):

\[3^2+6^2=45\]and \[7^2=49\] \[45\neq49\] In other words, the Pythagorean Theorem is not satisfied, and therefore it is not a right triangle.

OpenStudy (anonymous):

no~ the right triangle need to fit the formula a^2+b^2=c^2, and \[9+36\neq49\]